# Dual-Space Knowledge Distillation for Large Language Models

### v2

@@ 5.3 Representation Similarity between the Teacher and the Student
In the simulation experiment, we find that the current KD framework will lead to limited representation similarities between the student and the teacher (as shown in Figure 1(b)). Thus, we evaluate whether this phenomenon also holds in the real KD scenario. Specifically, we use cosine similarity and normalized inner product between output hidden states to represent the representation structure of a model. The detailed calculation of the structure distance is in Appendix E. Then we plot the average distance between the structure of the teacher and the student on 1000 training samples in Figure 3. It shows that on both types of representation structures, the current KD framework (Vanilla KD) only reduces minor distances between the teacher and the student compared to fine-tuning without KD (SFT). However, our DSKD achieves significantly lower distances between the teacher and the student, which indicates that DSKD can enhance the similarity between the student and the teacher.
@@ 6 Related Work
The white-box KD framework for language models stems from the standard KD method proposed by Hinton et al. (2015). As pre-trained language models (PLMs) become prevalent for various NLP tasks, numerous KD methods within this framework were proposed to compress the excessive model sizes of PLMs Sun et al. (2019); Sanh et al. (2019); Sun et al. (2020); Jiao et al. (2020). Besides minimizing the distance between distributions, there are also feature-based KD methods that distill the knowledge in intermediate hidden states and attention maps of the teacher model Jiao et al. (2020); Wang et al. (2020, 2021). Since LLMs are predominate for various tasks, several KD techniques have also been proposed for LLMs Gu et al. (2023); Ko et al. (2024); Wu et al. (2024); Xu et al. (2024). Unlike the previous work that follows the current white-box KD framework, we challenge this framework by revealing its inherent limitations and proposing a simple yet more effective and general KD framework as the solution.
@@
In the previous literature on KD, SimKD (Chen et al., 2022) was also proposed to share the teacher’s prediction head for KD, which was similar to the process of KD in the teacher space in our DSKD. However, the aim of SimKD is to equip the prediction head of the teacher model to the student model, and thus the student model will be larger after KD and suffer from higher inference costs. In contrast, our DSKD only leverages this process to transfer the representation information from the teacher and has no influence on the original model size of the student.
@@ A.2 Pseudo Code for Simulation Experiments
1class Teacher(nn.Module):
2 def __init__(self):
3 super(Teacher, self).__init__()
4 # the initial teacher hiddens are sampled from Gaussian Distribution N(0, 2)
5 self.hidden = torch.randn(100, 2) * 2
6 # the head contains 10000 classes
7 self.head = torch.randn(10000, 2)
8
9class Student(nn.Module):
10 def __init__(self):
11 super(Student, self).__init__()
12 # the initial student hiddens are sampled from Gaussian Distribution N(3, 1)
13 self.hidden = nn.Parameter(torch.randn(100, 2) + 3)
14 # the head contains 10000 classes
15 self.head = nn.Parameter(torch.randn(10000, 2))
16
17def kd_with_different_head(student, teacher):
18 student_logits = student.hidden.matmul(student.head.transpose(-1, -2))
19 # calculating logits with the respective heads
20 teacher_logits = teacher.hidden.matmul(teacher.head.transpose(-1, -2))
21 kd_loss = distance_func(student_logits, teacher_logits)
22 return kd_loss
23
24def kd_with_shared_head(student, teacher):
25 student_logits = student.hidden.matmul(student.head.transpose(-1, -2))
26 # calculating logits with the same head (student’s head)
27 teacher_logits = teacher.hidden.matmul(student.head.transpose(-1, -2))
28 kd_loss = distance_func(student_logits, teacher_logits)
29 return kd_loss

### v3

@@ 5.3 Representation Similarity between the Teacher and the Student
In the simulation experiment, we find that the current KD framework will lead to limited representation similarities between the student and the teacher (as shown in Figure 1(b)). Thus, we evaluate whether this phenomenon also holds in the real KD scenario. Since the dimensions are usually different for the teacher and student models, we measure the similarity of representation structures of the two models instead of their hidden states. Specifically, we use cosine similarity and normalized inner product between output hidden states to represent the representation structure of a model (see Eqn. (16) and (17) in Appendix E for the definitions). Then we calculate the L1 distance between the representation structures to reflect their similarity, where lower distance denotes higher similarity between representation structures (see Eqn. (18) and (19) in Appendix E for the detailed calculations). The average distances between the structure of the teacher and the student on 1000 training samples are plotted in Figure 3. It shows that on both types of representation structures, the current KD framework (Vanilla KD) only reduces minor distances between the teacher and the student compared to fine-tuning without KD (SFT). However, our DSKD achieves significantly lower distances between the teacher and the student, which indicates that DSKD can enhance the similarity between the student and the teacher.
@@ 6 Related Work
The white-box KD framework for language models stems from the standard KD method proposed by Hinton et al. (2015). As pre-trained language models (PLMs) become prevalent for various NLP tasks, numerous KD methods within this framework were proposed to compress the excessive model sizes of PLMs Sun et al. (2019); Sanh et al. (2019); Sun et al. (2020); Jiao et al. (2020). Besides minimizing the distance between distributions, there are also feature-based KD methods that distill the knowledge in intermediate hidden states and attention maps of the teacher model Jiao et al. (2020); Wang et al. (2020, 2021b). Additionally, white-box KD is also widely used in text generation tasks, such as neural machine translation Tan et al. (2019); Wang et al. (2021a); Zhang et al. (2023) and text summarization Chen et al. (2020); Liu et al. (2021). Since LLMs are predominate for various tasks, several KD techniques have also been proposed for LLMs Gu et al. (2023); Ko et al. (2024); Wu et al. (2024); Xu et al. (2024). Unlike the previous work that follows the current white-box KD framework, we challenge this framework by revealing its inherent limitations and proposing a simple yet more effective and general KD framework as the solution.
@@
In the previous literature on KD, SimKD (Chen et al., 2022) also proposed to share the teacher’s prediction head for KD, which was similar to the process of KD in the teacher space in our DSKD. However, the aim of SimKD is to equip the prediction head of the teacher model to the student model, and thus the student model will be larger after KD and suffer from higher inference costs. In contrast, our DSKD only leverages this process to transfer the representation information from the teacher and has no influence on the original model size of the student.
@@ 7 Conclusion
Acknowledgements
The research work described in this paper has been supported by the National Nature Science Foundation of China (No. 62476023, 61976016, 62376019, 61976015), and the authors would like to thank the anonymous reviewers for their valuable comments and suggestions to improve this paper.
@@ A.2 Pseudo Code for Simulation Experiments
class Teacher(nn.Module):
def __init__(self):
super(Teacher, self).__init__()
# the initial teacher hiddens are sampled from Gaussian Distribution N(0, 2)
self.hidden = torch.randn(100, 2) * 2
# the head contains 10000 classes
self.head = torch.randn(10000, 2)
class Student(nn.Module):
def __init__(self):
super(Student, self).__init__()
# the initial student hiddens are sampled from Gaussian Distribution N(3, 1)
self.hidden = nn.Parameter(torch.randn(100, 2) + 3)
# the head contains 10000 classes
self.head = nn.Parameter(torch.randn(10000, 2))
def kd_with_different_head(student, teacher):
student_logits = student.hidden.matmul(student.head.transpose(-1, -2))
# calculating logits with the respective heads
teacher_logits = teacher.hidden.matmul(teacher.head.transpose(-1, -2))
kd_loss = distance_func(student_logits, teacher_logits)
return kd_loss
def kd_with_shared_head(student, teacher):
student_logits = student.hidden.matmul(student.head.transpose(-1, -2))
# calculating logits with the same head (student’s head)
teacher_logits = teacher.hidden.matmul(student.head.transpose(-1, -2))
kd_loss = distance_func(student_logits, teacher_logits)
return kd_loss